# EvalMuse-40K: A Reliable and Fine-Grained Benchmark with Comprehensive Human Annotations for Text-to-Image Generation Model Evaluation

### v1

@@ 9 More Experiments
Table 6: Accuracy (%) of fine-grained evaluation method on different element categories. Align-Score represents the average fine-grained human annotation for each element category, indicating the degree of alignment with the image. 1,2,3 denote the use of Llava1.6 [28], mPLug-Owl3 [46], Qwen2-VL [40] for the VQA task respectively. And a./h. is the abbreviation for animal/human. All the above methods used a fixed-step (0.01) search to select the optimal threshold for binary classification, aiming to maximize overall accuracy.
@@ 9.2 More Qualitative Comparisons
As shown in Figs. 78, we present additional visual results to illustrate that our method provides improved alignment evaluation between generated images and prompts.
@@ 9.3 Structural Problem Evaluation
EvalMuse-40K’s annotation on structural problems allows us to measure the ability of MLLMs to identify structural problems in generated images. Such ability can be reflected by the model’s accuracy and recall. The accuracy can reflect the correctness and confidence of the model’s judgment, and recall can reflect the model’s ability to check all questions. We use the VQA method to ask the MLLM judgment questions, then, we apply softmax to the logit of the model outputs (“Yes.” and “No.”) to obtain the probability for each response.
Prompts. We use a unified template as “Does this generated image have any structural problems? Specifically, $<cls\_prompt>$? Please answer yes or no only.” where $<cls\_prompt>$ varies according to different tests and different structural problem labels. For the accuracy test, we only test on coarse-grained “animal”, “object” and “human-body” labels. For the recall test, we test all the annotated structural problems, including coarse-grained and fine-grained. The $<cls\_prompt>$ for testing are shown in Tab. 7.
Select Threshold. We define the threshold as the gate-value of the model answering “Yes.”, which means if the probability is larger than the threshold, the model’s output will be chosen “Yes.”.
Since MLLMs have different tendencies to answer Yes or No, if we simply use 0.5 as the threshold for answering Yes or No, it will lead to unfair testing. Therefore, before calculating the accuracy and recall, we first select an optimal threshold for each model.
@@
Accuracy. Considering that each model has different abilities to assess different categories of objects, we test all three coarse-grained categories for each image. The probability of model output “Yes.” is compared to the threshold to obtain the final output of the model. For each category, the accuracy rate is the number of correct answers divided by the total number of questions. The results of the accuracy test are shown in Tab. 8.
@@
Table 9: The recall rates of the three models. The recall rates of the three models are not high, which proves the ability of MLLMs to detect structural problems in images still requires improvement.
@@
Dreamina v2.0Pro [9] 3.74 0.821 0.793 0.706 0.747 0.689 0.756 0.700 0.580 0.662 0.747 0.477
DALLE 3 [32] 3.63 0.814 0.782 0.692 0.640 0.701 0.734 0.700 0.682 0.644 0.768 0.438
FLUX 1.1 [3] 3.47 0.819 0.758 0.660 0.642 0.638 0.686 0.673 0.607 0.596 0.671 0.362
Midjourney v6.1 [15] 3.33 0.807 0.736 0.637 0.693 0.625 0.619 0.718 0.659 0.599 0.716 0.285
SD 3 3.27 0.790 0.728 0.595 0.695 0.546 0.560 0.716 0.637 0.559 0.646 0.305
Playground v2.5 [22] 3.20 0.812 0.785 0.544 0.657 0.541 0.578 0.709 0.675 0.574 0.634 0.262
SDXL-Turbo [36] 3.15 0.788 0.714 0.494 0.659 0.487 0.567 0.671 0.665 0.551 0.644 0.306
HunyuanDiT [24] 3.08 0.794 0.753 0.555 0.666 0.524 0.576 0.682 0.705 0.586 0.648 0.247
Kandinsky3 [2] 3.08 0.793 0.723 0.541 0.652 0.513 0.583 0.681 0.661 0.564 0.665 0.291
SDXL [26] 2.99 0.786 0.717 0.467 0.623 0.463 0.533 0.677 0.660 0.531 0.607 0.276
PixArt-$\Sigma$ [4] 2.98 0.792 0.755 0.564 0.633 0.533 0.561 0.692 0.703 0.533 0.641 0.238
Kolors [38] 2.93 0.790 0.722 0.498 0.622 0.480 0.527 0.621 0.713 0.496 0.594 0.245
SDXL-Lightning [26] 2.93 0.788 0.729 0.478 0.619 0.458 0.534 0.619 0.600 0.528 0.609 0.274
SSD1B [29] 2.93 0.798 0.730 0.502 0.610 0.480 0.504 0.688 0.684 0.508 0.590 0.297
PixArt-$\alpha$ [6] 2.88 0.780 0.738 0.483 0.607 0.472 0.521 0.627 0.670 0.523 0.600 0.240
IF [8] 2.77 0.725 0.620 0.452 0.577 0.416 0.475 0.570 0.632 0.498 0.581 0.188
LCM-SDXL [29] 2.77 0.762 0.706 0.465 0.575 0.454 0.513 0.616 0.615 0.496 0.587 0.273
PixArt-$\delta$ [5] 2.73 0.768 0.718 0.455 0.565 0.432 0.486 0.634 0.685 0.496 0.574 0.207
LCM-SSD1B [29] 2.66 0.761 0.683 0.451 0.540 0.393 0.457 0.523 0.673 0.459 0.572 0.265
SD v2.1 [33] 2.42 0.698 0.590 0.354 0.502 0.363 0.431 0.532 0.559 0.398 0.528 0.190
SD v1.5 [33] 2.25 0.671 0.534 0.328 0.470 0.337 0.372 0.487 0.500 0.352 0.488 0.180
SD v1.2 [33] 2.25 0.659 0.515 0.315 0.471 0.377 0.393 0.498 0.547 0.349 0.493 0.181
@@ 10 T2I Model Alignment Evaluation
Benefiting from the strong correlation between FGA-BLIP2 and human preferences in image-text alignment evaluation, we employ it to assess and rank the performance of common T2I models in this task.

### v2

@@ 9 More Experiments
Table 6: Accuracy (%) of fine-grained evaluation methods on different element categories. Align-Score represents the average fine-grained human annotation for each element category, indicating the degree of alignment with the image. 1,2,3 denote the use of Llava1.6 [28], mPLug-Owl3 [46], Qwen2-VL [40] for the VQA task, respectively. And a./h. is the abbreviation for animal/human. All the above methods use a fixed-step (0.01) search to select the optimal threshold for binary classification, aiming to maximize overall accuracy.
@@ 9.2 More Qualitative Comparisons
As shown in Figs. 7 and 8, we present additional visual results to illustrate that our method provides improved alignment evaluation between generated images and prompts.
@@ 9.3 Structural Problem Evaluation
EvalMuse-40K’s annotation on structural problems allows us to measure the ability of MLLMs to identify structural problems in generated images. Such ability can be reflected by the model’s accuracy and recall. The accuracy can reflect the correctness and confidence of the model’s judgment, and recall can reflect the model’s ability to check all questions. We use the VQA method to pose judgment questions to the MLLM, then apply softmax to the logits of the model’s outputs (“Yes.” and “No.”) to obtain the probability for each response.
Prompts. We use a unified template as “Does this generated image have any structural problems? Specifically, $<cls\_prompt>$? Please answer yes or no only.” where $<cls\_prompt>$ varies according to different tests and different structural problem labels. For the accuracy test, we only test on coarse-grained “animal”, “object”, and “human-body” labels. For the recall test, we test all the annotated structural problems, including coarse-grained and fine-grained. The $<cls\_prompt>$ for testing are shown in Tab. 7.
Select Threshold. We define the threshold as the gate-value of the model answering “Yes.”, which means if the probability is larger than the threshold, the model’s output will be chosen “Yes.”. Since MLLMs have different tendencies to answer Yes or No, if we simply use 0.5 as the threshold for answering Yes or No, it will lead to unfair testing. Therefore, before calculating the accuracy and recall, we first select an optimal threshold for each model.
@@
Accuracy. Considering that each model has different abilities to assess different categories of objects, we test all three coarse-grained categories for each image. The probability of model output “Yes.” is compared to the threshold to obtain the final output of the model. For each category, the final accuracy rate is the average of accuracy of positive and negative samples. The results of the accuracy test are shown in Tab. 8.
@@
Table 9: The recall rates of the three models. The recall rates of the three models are relatively low, indicating that the ability of MLLMs to detect structural problems in images still requires significant improvement.
@@
Dreamina v2.0Pro [9] $3.74_{(1)}$ $0.821_{(1)}$ $0.793_{(1)}$ $0.706_{(1)}$ $0.747_{(1)}$ $0.689_{(2)}$ $0.756_{(1)}$ $0.700_{(4)}$ $0.580_{(19)}$ $0.662_{(1)}$ $0.747_{(2)}$ $0.477_{(1)}$
DALLE 3 [32] $3.63_{(2)}$ $0.814_{(3)}$ $0.782_{(3)}$ $0.692_{(2)}$ $0.732_{(2)}$ $0.701_{(1)}$ $0.734_{(2)}$ $0.700_{(4)}$ $0.682_{(6)}$ $0.644_{(2)}$ $0.768_{(1)}$ $0.438_{(2)}$
FLUX 1.1 [3] $3.47_{(3)}$ $0.819_{(2)}$ $0.758_{(4)}$ $0.660_{(3)}$ $0.694_{(4)}$ $0.638_{(3)}$ $0.686_{(3)}$ $0.673_{(11)}$ $0.607_{(17)}$ $0.596_{(4)}$ $0.671_{(4)}$ $0.362_{(3)}$
Midjourney v6.1 [15] $3.33_{(4)}$ $0.807_{(5)}$ $0.736_{(8)}$ $0.637_{(4)}$ $0.693_{(5)}$ $0.625_{(4)}$ $0.619_{(4)}$ $0.718_{(1)}$ $0.659_{(13)}$ $0.599_{(3)}$ $0.716_{(3)}$ $0.285_{(8)}$
SD 3 [10] $3.27_{(5)}$ $0.790_{(10)}$ $0.728_{(11)}$ $0.595_{(5)}$ $0.695_{(3)}$ $0.546_{(5)}$ $0.560_{(10)}$ $0.716_{(2)}$ $0.637_{(14)}$ $0.559_{(8)}$ $0.646_{(7)}$ $0.305_{(5)}$
Playground v2.5 [22] $3.20_{(6)}$ $0.812_{(4)}$ $0.785_{(2)}$ $0.544_{(8)}$ $0.657_{(8)}$ $0.541_{(6)}$ $0.578_{(6)}$ $0.709_{(3)}$ $0.675_{(7)}$ $0.574_{(6)}$ $0.634_{(10)}$ $0.262_{(13)}$
SDXL-Turbo [36] $3.15_{(7)}$ $0.788_{(12)}$ $0.714_{(16)}$ $0.494_{(12)}$ $0.659_{(7)}$ $0.487_{(10)}$ $0.567_{(8)}$ $0.671_{(12)}$ $0.665_{(10)}$ $0.551_{(9)}$ $0.644_{(8)}$ $0.306_{(4)}$
HunyuanDiT [24] $3.08_{(8)}$ $0.794_{(7)}$ $0.753_{(6)}$ $0.555_{(7)}$ $0.666_{(6)}$ $0.524_{(8)}$ $0.576_{(7)}$ $0.682_{(8)}$ $0.705_{(2)}$ $0.586_{(5)}$ $0.648_{(6)}$ $0.247_{(14)}$
Kandinsky3 [2] $3.08_{(8)}$ $0.793_{(8)}$ $0.723_{(12)}$ $0.541_{(9)}$ $0.652_{(9)}$ $0.513_{(9)}$ $0.583_{(5)}$ $0.681_{(9)}$ $0.661_{(11)}$ $0.564_{(7)}$ $0.665_{(5)}$ $0.291_{(7)}$
SDXL [26] $2.99_{(10)}$ $0.786_{(14)}$ $0.717_{(15)}$ $0.467_{(15)}$ $0.623_{(11)}$ $0.463_{(14)}$ $0.533_{(12)}$ $0.677_{(10)}$ $0.660_{(12)}$ $0.531_{(11)}$ $0.607_{(12)}$ $0.276_{(9)}$
PixArt-$\Sigma$ [4] $2.98_{(11)}$ $0.792_{(9)}$ $0.755_{(5)}$ $0.564_{(6)}$ $0.633_{(10)}$ $0.533_{(7)}$ $0.561_{(9)}$ $0.692_{(6)}$ $0.703_{(3)}$ $0.533_{(10)}$ $0.641_{(9)}$ $0.238_{(17)}$
Kolors [38] $2.93_{(12)}$ $0.790_{(10)}$ $0.722_{(13)}$ $0.498_{(11)}$ $0.622_{(12)}$ $0.480_{(11)}$ $0.527_{(13)}$ $0.621_{(15)}$ $0.713_{(1)}$ $0.496_{(16)}$ $0.594_{(14)}$ $0.245_{(15)}$
SDXL-Lightning [26] $2.93_{(12)}$ $0.788_{(12)}$ $0.729_{(10)}$ $0.478_{(14)}$ $0.619_{(13)}$ $0.458_{(15)}$ $0.534_{(11)}$ $0.619_{(16)}$ $0.600_{(18)}$ $0.528_{(12)}$ $0.609_{(11)}$ $0.274_{(10)}$
SSD1B [29] $2.93_{(12)}$ $0.798_{(6)}$ $0.730_{(9)}$ $0.502_{(10)}$ $0.610_{(14)}$ $0.480_{(11)}$ $0.504_{(16)}$ $0.688_{(7)}$ $0.684_{(5)}$ $0.508_{(14)}$ $0.590_{(15)}$ $0.297_{(6)}$
PixArt-$\alpha$ [6] $2.88_{(15)}$ $0.780_{(15)}$ $0.738_{(7)}$ $0.483_{(13)}$ $0.607_{(15)}$ $0.472_{(13)}$ $0.521_{(14)}$ $0.627_{(14)}$ $0.670_{(9)}$ $0.523_{(13)}$ $0.600_{(13)}$ $0.240_{(16)}$
IF [8] $2.77_{(16)}$ $0.725_{(19)}$ $0.620_{(19)}$ $0.452_{(18)}$ $0.577_{(16)}$ $0.416_{(18)}$ $0.475_{(18)}$ $0.570_{(18)}$ $0.632_{(15)}$ $0.498_{(15)}$ $0.581_{(17)}$ $0.188_{(20)}$
LCM-SDXL [29] $2.77_{(16)}$ $0.762_{(17)}$ $0.706_{(17)}$ $0.465_{(16)}$ $0.575_{(17)}$ $0.454_{(16)}$ $0.513_{(15)}$ $0.616_{(17)}$ $0.615_{(16)}$ $0.496_{(16)}$ $0.587_{(16)}$ $0.273_{(11)}$
PixArt-$\delta$ [5] $2.73_{(18)}$ $0.768_{(16)}$ $0.718_{(14)}$ $0.455_{(17)}$ $0.565_{(18)}$ $0.432_{(17)}$ $0.486_{(17)}$ $0.634_{(13)}$ $0.685_{(4)}$ $0.496_{(16)}$ $0.574_{(18)}$ $0.207_{(18)}$
LCM-SSD1B [29] $2.66_{(19)}$ $0.761_{(18)}$ $0.683_{(18)}$ $0.451_{(19)}$ $0.540_{(19)}$ $0.393_{(19)}$ $0.457_{(19)}$ $0.523_{(20)}$ $0.673_{(8)}$ $0.459_{(19)}$ $0.572_{(19)}$ $0.265_{(12)}$
SD v2.1 [33] $2.42_{(20)}$ $0.698_{(20)}$ $0.590_{(20)}$ $0.354_{(20)}$ $0.502_{(20)}$ $0.363_{(21)}$ $0.431_{(20)}$ $0.532_{(19)}$ $0.559_{(20)}$ $0.398_{(20)}$ $0.528_{(20)}$ $0.190_{(19)}$
SD v1.5 [33] $2.25_{(21)}$ $0.671_{(21)}$ $0.534_{(21)}$ $0.328_{(21)}$ $0.470_{(22)}$ $0.337_{(22)}$ $0.372_{(22)}$ $0.487_{(22)}$ $0.500_{(22)}$ $0.352_{(21)}$ $0.488_{(22)}$ $0.180_{(22)}$
SD v1.2 [33] $2.25_{(21)}$ $0.659_{(22)}$ $0.515_{(22)}$ $0.315_{(22)}$ $0.471_{(21)}$ $0.377_{(20)}$ $0.393_{(21)}$ $0.498_{(21)}$ $0.547_{(21)}$ $0.349_{(22)}$ $0.493_{(21)}$ $0.181_{(21)}$
@@ 10 T2I Model Alignment Evaluation
Benefiting from the strong correlation between FGA-BLIP2 and human preferences in image-text alignment evaluation, we employ it to assess and rank the performance of widely used T2I models.